OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 7, b = 7, c = 5 A. A = 70°, B = 70°, C = 40° B. A = 69°, B = 69°, C = 42° C. A = 42°, B = 69°, C = 69° D. A = 69°, B = 42°, C = 69°

OpenStudy (anonymous):

hey

OpenStudy (anonymous):



OpenStudy (anonymous):

law of cosines for angle C: c^2 = a^2 + b^2 - 2abcosC 25 = 49 + 49 - 2 * 49 * cosC -> cosC = 0.7448979... C = 41.85 -> C = 42 degrees now for A and B : A+B+C =180 A+B = 180 - 42 A+B = 138 since A = B we can find them : 2A = 138 A = B = 69 ANSWER IS B :)

OpenStudy (anonymous):

thank u !!!